hello I am daisy and im having trouble how to translate words into equation form. I would be glad if someone could help me . I've tried solving it a couple of times , but i seem to get the answer wrong .
Problem : The perimeter of a rectangle is 186 feet and the length is 75 feet more than the width . What are the dimensions of the rectangle ?

This is how i did the problem :
x= the dimensions 75=
equation : l*w
2(75+w)+2w=186
150+w+2w=186
75+(12)=87
The length would equal 87 ?

150+3w=186
-150 -150
3w = 36
then i would divide by 3
so that would give me 12 for the width .
Would that be the right answer ?



Answer by mananth(16946)

2(w+75)+2w=186
2w+150+2w=186
4w+150=186
4w=36
w=9